OpenStudy (anonymous):

roots for x^3-6x^2+7x-2

OpenStudy (anonymous):

x1=1, x2=(5-sqrt17)/2 = 0.438447, x3=(5+sqrt17)/2 = 4.56155

OpenStudy (anonymous):

x^3-6x^2+7x-2=0 =[x-1][x-(5-sqrt17)/2][x-(5+sqrt17)/2]=0 x1=1, x2=(5-sqrt17)/2 = 0.438447, x3=(5+sqrt17)/2 = 4.56155 answer

OpenStudy (anonymous):

here i used by trial and error, (x^3-6x^2+7x-2)/(x-1)=x^2 -5x +2 nowyou can use the quadratic formula to get the other two answer